John has 50 comic books and he increases his collection by 3 books per week


1
Expert's answer
2021-11-29T19:33:41-0500

Let "x=" the number of weeks. Then the John's collection "y" will be "y=50+3x" books after "x" weeks.


"1\ month=4\ weeks"

"y(4)=50+3(4)=62"

The John's collection will be "62" books after one month.